# Eddeep: Fast eddy-current distortion correction for diffusion MRI with deep learning

### v2

@@ 1 Introduction
Recently, major advances in image processing have been enabled by deep-learning (DL). DL-based registration methods [15, 13] demonstrated state-of-the-art alignment accuracy, doing so in a split second at inference. They have already been used for the correction of susceptibility distortions [22, 23], but not yet for eddy-current ones because the issue of disrupted correspondences holds. But DL has also revolutionized the domain of image translation where conditional GAN (cGAN) [11] architectures like pix2pix [9] have been used in medical imaging for T1-weighted to $b=0$ image translation in [12].
@@ 4.3 Implementation details
The code is available at: github.com/CIG-UCL/eddeep. It uses tools from VoxelMorph111github.com/voxelmorph/voxelmorph and SimpleITK222simpleitk.org.
@@ 4.4 Results
Registration model: Those models were compared against FSL Eddy, the state-of-the-art tool for this task. We used, for each b-value, the voxel-wise intensity standard deviation across volumes near the interface between cortical gray matter and CSF as an indicator of the amount of misalignment. This is motivated by the observation that mismatch of this interface across volumes will generate higher values of standard deviation compared to when this interface is well aligned across volumes. To quantify this, we extracted a mask of this interface (subtraction of a dilated (2 iterations) and eroded (iterations) brain mask from FSL BET [20]). Results are shown in Fig. 3. Correction is better for full models where the training of the registration was done with a sample of size M. Among those, interestingly, $R_{M}^{+}\circ T_{S}$ does not seem to be penalized by the training sample size S of its translator. Correction with the best performing models lead to output very similar to FSL Eddy.

### v3

@@ 1 Introduction
Recently, new approaches to geometric distortion corrections have been developed, taking advantage of major advances in image processing enabled by deep-learning (DL). In particular DL-based image registration methods [15, 13] demonstrated state-of-the-art alignment accuracy, doing so in a split second at inference. These have already successfully been applied to the correction of susceptibility distortions [22, 23], but not, to the best of our knowledge, to eddy distortions, likely because of the difficulty to handle disrupted correspondences. We hypothesize that this difficulty may be addressed by leveraging similar advances in image translation enabled by DL where conditional GAN (cGAN) [11] architectures like pix2pix [9] have been used in medical imaging for T1-weighted to $b=0$ image translation in [12].
@@ 4.3 Implementation details
The code is available at: https://github.com/CIG-UCL/eddeep. It uses tools from VoxelMorph [^1] https://github.com/voxelmorph/voxelmorph and SimpleITK [^2] https://simpleitk.org.
@@ 4.4 Results
Registration model: Those models were compared against FSL Eddy, the state-of-the-art tool for this task. We used, for each b-value, the voxel-wise intensity standard deviation across volumes near the interface between cortical gray matter and CSF as an indicator of the amount of misalignment. This is motivated by the observation that mismatch of this interface across volumes will generate higher values of standard deviation compared to when this interface is well aligned across volumes. To quantify this, we extracted a mask of this interface (subtraction of a dilated (2 iterations) and eroded (4 iterations) brain mask from FSL BET [20]). Results are shown in Fig. 3. Correction is better for full models where the training of the registration was done with a sample of size M. Among those, interestingly, $R_{M}^{+}\circ T_{S}$ does not seem to be penalized by the training sample size S of its translator. Correction with the best performing models lead to output very similar to FSL Eddy.
@@ 5 Discussion and conclusion
{credits}
5.0.1 Acknowledgements
This work is funded by the Medical Research Council (MRC): (MR/T046473/1; AL, HZ), Innovate UK (No. 10036158; HA, AL, HZ), Le Fonds de la Recherche Scientifique (FRS-FNRS) and the Research Foundation Flanders (FWO) as part of Excellence Of Science (EOS) project Memodyn (No. 30446199; AL, HZ, PP) and FRS-FNRS (WS).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.